OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2?

OpenStudy (jdoe0001):

same way as the previous one

OpenStudy (jdoe0001):

use the point-slope form and plug in the values

OpenStudy (anonymous):

y= 2x+5

OpenStudy (anonymous):

so y-5=2(x-3) ?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

so y-5=2(x-3) \(\Large \checkmark\)

OpenStudy (anonymous):

thank you :)